HETAV PATEL

Grade 11,

Hi... I am Vivek ,studying in 12th standard.I have a question here !!!! Apiece of equipment cost a certain factory Rs. 600,000. If it depreciates in value, 15% the first year, 13.5 % the next year, 12% the third year, and so on, what will be its value at the end of 10 years, all percentages applying to the original cost?

Hi... I am Vivek ,studying in 12th standard.I have a question here !!!!


Apiece of equipment cost a certain factory Rs. 600,000. If it depreciates in value, 15% the first year, 13.5 % the next year, 12% the third year, and so on, what will be its value at the end of 10 years, all percentages applying to the original cost?

Grade:

2 Answers

raja kapoor
15 years ago

Let the cost of an equipment is Rs. 100.
Now the percentages of depreciation at the end of 1st,2nd,3rd years are 15, 13.5,12, which are in A.P., with a=15 and d= -1.5.

Hence, percentage of depreciation in the tenth year = a + (10-1) d = 15 + 9 (-1.5) = 1.5
Also total value depreciated in 10 years = 15 + 13.5 + 12 + ... + 1.5 = 82.5

Hence, the value of equipment at the end of 10 years=100 - 82.5 = 17.5.

The total cost being Rs. 6,00,000/100 * 17.5 = Rs. 1,05,000.

I think my answer is correct

shruti seth
15 years ago

ya ....Ur answer is Correct !!!!!!!